# Auditing Political Exposure Bias: Algorithmic Amplification on Twitter/$\mathbb{X}$ Approaching the 2024 U.S. Presidential Election

### v2

@@ Methods
The parameters of the exponential decay function are informed by findings from studies on platforms like TikTok and YouTube (Guinaudeau et al., 2022), which indicate that the top 20% of an account’s videos receive more than 70% of the views. Using this as a reference, we assume that the top 20% of tweets in a timeline similarly capture the majority of user attention, and we calibrate our decay model accordingly. For instance, for a neutral account with an average timeline length of 500, the exponential decay function is defined as:
@@
To measure whether exposure is evenly distributed among users or dominated by a few accounts, we employ the Gini coefficient, a widely used measure to quantify inequality (Bartley et al., 2023). The Gini coefficient ranges from 0 to 1, where 0 indicates perfect equality (all users have the same exposure) and 1 signifies maximum inequality (exposure is concentrated among a few accounts). In our specific case, the Gini coefficient $G$ is calculated as:
@@
where $E_{i}$ and $E_{j}$ represent the exposure metrics—weighted occurrence per 1,000 tweets—of users $i$ and $j$ in a monitoring account’s timeline, $n$ is the total number of users, and $\bar{E}$ is the mean exposure metric across all users. A higher Gini coefficient indicates greater inequality in exposure distribution, suggesting that a small number of users dominate exposure in the timeline, while a lower coefficient suggests a more even distribution among users.
Amplification Measure
@@
Here, $V_{\text{left}}$ is the set of left-leaning accounts, and $V_{\text{balanced}}$ is the set of balanced accounts. $E_{v,u}$ denotes the weighted occurrence per 1,000 tweets of account $u$ in the timelines of account $v$. This amplification ratio quantifies the extent to which a user’s exposure is increased or decreased when viewed by left-leaning monitoring accounts compared to the balanced baseline. The calculation for right-leaning monitoring accounts follows a similar approach.
About the Team
The 2024 Election Integrity Initiative is led by Emilio Ferrara and Luca Luceri and carried out by a collective of USC students and volunteers whose contributions are instrumental to enable these studies. Special thanks to Marco Minici and Federico Cinus for comments on this study. The authors are also grateful to the following HUMANS Lab’s members for their tireless efforts on this project: Ashwin Balasubramanian, Leonardo Blas, Charles ‘Duke’ Bickham, Keith Burghardt, Sneha Chawan, Vishal Reddy Chintham, Eun Cheol Choi, Srilatha Dama, Priyanka Dey, Isabel Epistelomogi, Saborni Kundu, Grace Li, Richard Peng, Gabriela Pinto, Jinhu Qi, Ameen Qureshi, Namratha Sairam, Tanishq Salkar, Srivarshan Selvaraj, Kashish Atit Shah, Gokulraj Varatharajan, Reuben Varghese, Jinyi Ye, Siyi Zhou, and Vito Zou. Previous memos: (Ferrara, 2024a, b; Pinto et al., 2024; Minici et al., 2024; Blas et al., 2024; Balasubramanian et al., 2024; Cinus et al., 2024; Shah et al., 2024)
Appendix
Top Recommended Users in Left-Leaning, Right-Leaning, and Balanced Accounts
@@
Figure 5, figure 6 and figure 7 display the top 20 recommended users in left-leaning, right-leaning, and balanced accounts, ranked by their average weighted occurrence per 1,000 tweets. Each box in the boxplot represents the distribution of exposure across all accounts in each group, with red indicating right-leaning users and blue indicating left-leaning users. Political leanings of users are inferred based on publicly available data, which may be subject to inaccuracies or changes over time.
@@
Profile Information of the Top Recommended Users Across Account Groups
@@ 11 AMAZlNGNATURE Nature is Amazing Animals Nature Discovery 4496570
21 ProjectLincoln The Lincoln Project “You cannot escape the responsibility of tomorrow by evading it today.” – Abraham Lincoln | Home of #TheBreakdown and LP Podcast 2994780
@@ 28 DC_Draino DC_Draino Rogan O’Handley 1855070
32 PopBase Pop Base Pop Base is your best source for all pop culture related entertainment, news, award show coverage, chart updates, statistics and more. | email@popbase.tv 1683990
33 CollinRugg Collin Rugg Co-Owner of Trending Politics | Investor | American 1561596
@@
40 Riley_Gaines_ Riley Gaines Host of Gaines for Girls podcast | Author of Swimming Against the current | TPUSA contributor | Director of the Riley Gaines Center 1283591
@@ 49 Dexerto Dexerto The leading source for influencer, streamer, gaming, and viral content 980351
56 Angry_Staffer Angry Staffer Not a WH Staffer | Politics, NatSec, and Snark - Your Mileage May Vary | Subscribe to my Patreon newsletter for free: https://t.co/Kj4zTIcPyk | 609103
@@
60 acnewsitics Alex Cole Software Engineer & Pilot | Progressive Follow @newsitics & https://t.co/Retehye9rD 286475
@@ 61 Logically_JC John Collins Dad Husband Low-Key Nerd EdD / JD 225758
66 canlesofficial Canles Engineered for walking | Comfy & versatile footwear for life’s adventures | Breathable, lightweight designs 7407

### v3

@@ 3.3. Exposure Evaluation Metric
The parameters of the exponential decay function are informed by findings from studies on platforms like TikTok and YouTube (Guinaudeau et al., 2022), which indicate that the top 20% of an account’s videos receive more than 70% of the views. Using this as a reference, we assume that the top 20% of tweets in a timeline similarly capture the majority (70%) of user attention, and we calibrate our decay model accordingly. For instance, for a neutral account with an average timeline length of 500, the exponential decay function is defined as:
@@ 3.4. Inequality Measure
To measure whether exposure is evenly distributed among users or dominated by a few accounts, we employ the Gini coefficient, a widely used measure to quantify inequality (Bartley et al., 2023; Farris, 2010). The Gini coefficient ranges from 0 to 1, where 0 indicates perfect equality (all users have the same exposure) and 1 signifies maximum inequality (exposure is concentrated among a few accounts). In our specific case, the Gini coefficient $G$ is calculated as:
@@
where $E_{i}$ and $E_{j}$ represent the exposure metrics—weighted occurrence per 1,000 tweets—of users $i$ and $j$ in a monitoring account’s timeline, $n$ is the total number of users, and $\bar{E}$ is the mean exposure metric across all users. A higher Gini coefficient indicates greater inequality in exposure distribution, suggesting that a small number of users dominate exposure in the timeline, while a lower coefficient suggests a more even distribution among users. To complement this analysis, we use the Lorenz curve (Gastwirth, 1971) as a visual representation of exposure inequality.
3.5. Amplification Measure
@@
Here, $V_{\text{left}}$ is the set of left-leaning accounts, and $V_{\text{balanced}}$ is the set of balanced accounts. $E_{v,u}$ denotes the weighted occurrence per 1,000 tweets of account $u$ in the timelines of account $v$. This amplification ratio quantifies the extent to which a user’s exposure is increased or decreased when viewed by left-leaning monitoring accounts compared to the balanced baseline. A positive mean amplification ratio indicates amplification, while a negative ratio indicates de-amplification. The calculation for right-leaning monitoring accounts follows a similar approach.
4. Results
4.1. Out-of-Network Exposure Inequality Among Different Political Profiles (RQ1)
RQ1 explores the extent to which personalized recommendations in $\mathbb{X}$ exhibit exposure inequality among users and how these inequalities vary between partisan accounts. To address this question, we use the Gini coefficient, a standard measure of inequality that quantifies disparities in exposure by calculating how concentrated exposure is across a set of users. Detailed descriptions of the Gini coefficient calculation and the exposure metric are provided in the Methods section. For each sock-puppet monitoring account, we compute its Gini coefficient with respect to all recommended users in that account’s timelines.
Figure 1. Distribution of Gini coefficient across different groups of accounts. Significant disparities are found in all pairwise comparisons (Mann-Whitney U test: $p$ 0.001), with right-leaning users experiencing the highest out-of-network exposure inequality.
[…]
Gini coefficient violin plot
Figure 2. Lorenz curves for different groups of accounts. Each curve represents the average Lorenz curve for all accounts in the group, with error bars indicating the standard deviations at each cumulative point.
[…]
Lorenz curve cumulative plot
Figure 1 presents the distribution of Gini coefficients across different account groups: Left-Leaning, Right-Leaning, Balanced, and Neutral. The average Gini coefficient across all groups exceeds 0.45, which suggests a moderate to high level of inequality in exposure on the $\mathbb{X}$ platform. Similarly, as shown in Figure 2, the Lorenz curves for all account groups deviate substantially from the line of equality (dashed black line). The greater the curvature of the Lorenz curve, the higher the inequality in exposure. This indicates that algorithmic exposure is concentrated among certain users rather than evenly distributed.
@@ 4.1. Out-of-Network Exposure Inequality Among Different Political Profiles (RQ1)
Figure 3. Top 20 recommended users for neutral accounts, ranked by their average weighted occurrence per 1,000 tweets. Each box in the boxplot shows the distribution of exposure across all neutral accounts, with red and blue colors indicating right- and left-leaning users, respectively. The figure suggests that right-leaning users are more frequently recommended than left-leaning users in the algorithm’s out-of-network recommendations for neutral accounts.
[…]
Top 20 recommendations neutral
Notably, right-leaning users experience the highest exposure inequality, followed by balanced and left-leaning users. The Mann-Whitney U test reveals that the differences in Gini coefficients between all pairs of groups are significant at the 0.001 level, underscoring meaningful disparities in exposure inequality across these groups. This suggests that the algorithm’s out-of-network tweet recommendations for right-leaning users are more centralized, reflecting a stronger popularity bias, where a few users dominate exposure. In contrast, neutral users—who do not follow anyone—receive the most diverse recommendations, potentially due to algorithmic cold start, i.e., the absence of information about user preferences that typically informs recommendations (Yuan and Hernandez, 2023).
Now that we understand that out-of-network exposures are skewed toward certain users, an important question arises: Who are these users? Here, we are particularly interested in neutral accounts, which provide an unbiased look at the algorithm’s default behavior. Since neutral accounts are critical for detecting bias, we took particular care in their setup to ensure neutrality. Neutral accounts follow no other accounts and, therefore, receive exclusively out-of-network recommendations. This configuration limits any bias that could arise from following choices, aiming to capture a baseline view of how the algorithm behaves when no user preferences are specified. However, it is worth noting that certain factors, such as $\mathbb{X}$’s default settings or trending topics, could still introduce slight biases into these recommendations.
Figure 3 displays the top 20 recommended users for neutral accounts, ranked by their weighted occurrence per 1,000 tweets. Each box in the boxplot represents the distribution of this exposure metric across all neutral accounts. Boxes are colored red or blue to indicate whether the user is right- or left-leaning, based on publicly available data, including $\mathbb{X}$ user profile descriptions and external sources such as Wikipedia. A user’s political stance is classified as left- or right-leaning if they are affiliated with a political party or a media outlet with a recognized ideological alignment. A qualitative inspection reveals that right-leaning users appear more frequently among the top recommendations than left-leaning users. To quantify this difference, we use the “weighted occurrences per 1,000 tweets” metric: among the top 20 recommended users, right-leaning users make up 30.16% of exposure, compared to 12.92% for left-leaning users. This disparity persists as we expand the pool, with right-leaning users making up 35.26% of exposure in the top 50 (versus 22.34% for left-leaning users) and 31.39% in the top 100 (versus 20.83% for left-leaning users).
Notably, balanced accounts receive a roughly even mix of left- and right-leaning recommendations, whereas left- and right-leaning accounts predominantly receive recommendations from ideologically aligned users. In the Appendix, interested readers can find the top 20 recommendations for left-leaning, right-leaning, and balanced account groups, highlighting the most amplified users within each account category. A detailed table describing these users’ public information is also provided in the Appendix.
4.2. Differential (De-)Amplification of Political Content Among Partisan Accounts (RQ2)
[…]
amplified_accounts_left_50
[…]
amplified_accounts_right_50
Figure 4. Amplification ratio of the top 50 recommended users in left-leaning (top) and right-leaning (bottom) accounts, compared to the baseline of balanced accounts. Colored bars indicate a significant difference in exposure metrics (weighted occurrence per 1,000 tweets) between the groups at the 0.05 significance level (using the Mann-Whitney U test), while gray bars indicate no significant difference. Usernames are displayed in blue (left-leaning) or red (right-leaning) based on their political stance, according to publicly available data.
To address RQ2 and evaluate the amplification of certain users in partisan accounts’ timelines, we introduce the “mean amplification ratio” metric inspired by Huszár et al. (2022), as detailed in the Methods section. Figure 4 shows the amplification ratio of the top 50 recommended users in left-leaning and right-leaning accounts, compared to a baseline observed in politically balanced accounts’ timelines. Colored bars indicate a significant difference in exposure metrics (weighted occurrence per 1,000 tweets) between groups at the 0.05 significance level (using the Mann-Whitney U test), while gray bars indicate no significant difference.
A qualitative inspection reveals that left-leaning sock-puppet accounts tend to see left-leaning users amplified, and right-leaning users de-amplified, with the opposite pattern observed for right-leaning accounts. For instance, in left-leaning accounts, the top three amplified users are Ron Filipkowski (a former federal prosecutor known for his criticisms of conservative figures), Mueller, She Wrote (a political commentary and investigative journalism account with a liberal stance), and George Takei (an American actor, author and Democrat activist). In contrast, the most de-amplified accounts are Elon Musk (CEO of Twitter/$\mathbb{X}$, who has recently shared conservative viewpoints), Charlie Kirk (a conservative political activist), and Jack Posobiec (a right-wing media personality and political activist). This suggests that left-leaning timelines prioritize left-aligned figures while downplaying right-leaning accounts.
On the contrary, for right-leaning accounts, the top three accounts with the highest amplification in right-leaning timelines are catturd2 (a right-wing influencer known for political satire), atensnut (a conservative commentator), and DC_Draino (Rogan O’Handley, a right-wing political commentator). Conversely, the most de-amplified accounts are JoJoFromJerz (a left-leaning political influencer), acnewsitics (a liberal-leaning news commentator), and Tristan Snell (a pro-democrat lawyer and legal commentator). This pattern highlights the algorithm’s tendency to amplify conservative figures more heavily in right-leaning timelines while reducing exposure to left-leaning accounts.
To further illustrate this trend, usernames are displayed in blue (left-leaning) or red (right-leaning) based on their political stance, which is inferred from publicly available data (may be subject to inaccuracies or changes over time). As shown in Figure 4, top liberal and conservative voices are amplified more than 50% above baseline for left- and right-leaning users, respectively. Given that our sock-puppet accounts only follow a few moderately partisan media and politicians, it suggests that once a new user begins following a few partisan accounts, their algorithmic recommendations quickly become filled with like-minded voices.
Interestingly, we observe that amplified users in left-leaning group experience a slightly higher magnitude of amplification compared to those in right-leaning group ($M_{left}=36.76\%$, $M_{right}=30.29\%$, Mann-Whitney U $p<0.05$). However, there are no significant differences in the extent of de-amplification between the two groups.
5. Discussion & Conclusions
In this study, we present a six-week audit of algorithmic recommendations on $\mathbb{X}$’s “For You” timelines during the course of the 2024 U.S. Election. Using 120 sock-puppet accounts with left-leaning, right-leaning, balanced, and neutral political orientations, we observe that $\mathbb{X}$ skews exposure toward a select few high-popularity accounts for all users, with right-leaning users experiencing the highest level of inequality. Both left- and right-leaning users see amplified exposure to accounts aligned with their political stance, while exposure to opposing viewpoints is reduced. Additionally, analysis of neutral accounts with no follow activity reveals a default right-leaning bias in the platform’s recommendations.
Our analysis of exposure inequality aligns with previous studies on algorithmic bias, which have reported similar amplification patterns within users’ in-network content (Bartley et al., 2023). However, our findings diverge from earlier research suggesting that personalized recommendations tend to be more centrist in political stance (Chen et al., 2020; Bouchaud et al., 2023; Wang et al., 2024). This discrepancy perhaps highlights a shift in $\mathbb{X}$’s algorithmic behavior, which might have moved away from promoting moderate content to reinforcing users’ existing preferences more explicitly, especially in out-of-network recommendations. The results also add to the growing body of literature indicating that right-leaning accounts are often more prominently featured in algorithmic curation (Graham and Andrejevic, 2024), a trend seen here in the default bias toward right-leaning content for new or neutral accounts.
Another noteworthy observation is that, unlike prior research, which has primarily examined the amplification of tweets from media outlets (Wang et al., 2024) and political figures—especially elected legislators from major political parties (Huszár et al., 2022)—our findings reveal that $\mathbb{X}$’s algorithm now also amplifies political commentators and influencers. This trend is most pronounced in the recommendations for neutral accounts, suggesting a shift in the algorithm’s prioritization toward these types of voices. This shift could be influenced by recent claims that $\mathbb{X}$ prioritizes verified and paid subscription accounts [^2] Tweet from Twitter/X CEO Elon Musk
https://x.com/elonmusk/status/1650731557164818437?lang=en
, potentially amplifying influencers who invest in these platform features. The prominence of these non-institutional voices in political content raises questions about the influence of individual commentators on public opinion, as their perspectives may carry a more personal or sensational tone compared to traditional media sources. Adding to the concerns, recent investigations uncovered state-sponsored foreign interference operations with financial backing of prominent political influencers. [^3] Justice Department Disrupts Covert Russian Government-Sponsored Foreign Malign Influence Operation Targeting Audiences in the United States and Elsewhere https://www.justice.gov/opa/pr/justice-department-disrupts-covert-russian-government-sponsored-foreign-malign-influence This underscores the need for further examination into how the recommendation algorithm’s priorities may shape political engagement and public discourse, especially during critical periods like an election year.
Implications and Future Research.
Our research findings offer both theoretical and practical implications regarding the algorithm’s influence on echo chambers and the design of transparency-aware content recommendation algorithms. The $\mathbb{X}$ algorithm’s amplification of ideologically aligned out-of-network accounts, along with the reduced exposure to opposing viewpoints, suggests that algorithmic recommendations can reinforce echo chambers not just in the composition of social networks (Duskin et al., 2024) but also in the ideological framing of content circulating in the network. The increased prominence of non-institutional voices, such as verified political commentators and influencers, further exacerbates this issue by potentially introducing sensationalism and misinformation into these echo chambers (de Arruda et al., 2024). Additionally, the default right-leaning bias observed for neutral accounts suggests that new users are likely to encounter partisan content early in their engagement with the platform. This raises concerns about how early algorithmic shaping of timelines might influence political perspectives and preferences. Future research could address these concerns by 1) systematically comparing in-network and out-of-network exposure biases and 2) conducting user studies to investigate how algorithmically curated timelines influence political attitudes over time (see “sociotechnical audit” (Wang et al., 2024)).
The study also provides practical considerations for designing fair and transparent algorithms. Current recommendation systems appear to disproportionately amplify high-popularity accounts, creating inequality in exposure that may result in less personalized and miscalibrated recommendations for certain user groups (Abdollahpouri et al., 2020). Fairness algorithms could address this by factoring in diversity constraints that balance the exposure of popular and less popular accounts. Platforms should enhance transparency around how algorithms prioritize specific users, particularly verified and paid subscription accounts for the Twitter/$\mathbb{X}$ scenario. Future research should focus on monitoring algorithmic shifts and developing transparency standards during high-stakes periods such as elections, public health crises, and social unrest, where equitable and informed public discourse is critical.
Limitations.
We acknowledge several limitations in our research. First, the study is conducted during a six-week period leading up to the 2024 U.S. elections, a politically charged time that may differ from other contexts. This temporal limitation could affect the reproducibility of our results in less politically sensitive periods. Second, we deliberately avoid inducing interactions between sock-puppet accounts and algorithmic recommendations to isolate baseline biases in the recommendation system. Although our sock-puppet auditing method ensures precise control over account behaviors, it does not account for personalization or the dynamics of user activity. Since the sock-puppet accounts do not engage with tweets (e.g., clicking, responding, or retweeting), our study does not capture the effects of user-algorithm interactions on political exposure bias. Third, potential confounding factors, such as pre-selected interests, age, and location, may influence algorithmic recommendations for neutral accounts. While we took care to randomize these settings, their residual effects cannot be entirely ruled out. Forth, the use of balanced accounts as a baseline for measuring exposure biases may not fully capture the platform’s broader algorithmic behavior across diverse user demographics or global political contexts, potentially limiting the generalizability of our findings.
Ethical Statement.
Throughout our research process, we have adhered to stringent ethical standards to ensure the integrity and societal responsibility of our work. Our sock-puppet accounts were designed solely to follow media and public figures, observe, and collect data, without engaging in any interactions with real users on the $\mathbb{X}$ platform, thereby avoiding disruptions to other users’ experiences. All personal-identifiable information utilized in this study pertains exclusively to public figures and is derived from publicly available data. Additionally, we have carefully considered the societal impacts of our research. To mitigate risks of overgeneralization or misinterpretation, we provide thorough contextual information and openly address the limitations of our findings. While acknowledging these potential risks, we posit that our work could contribute to the development of algorithmic transparency standards and inform platform responsibilities during politically sensitive periods in the long term.
Appendix A Data Collection Details
Figure 5 display the number of active accounts and the total tweets collected daily. Data collection for neutral monitoring accounts began around October 2, 2024, and reached a stable deployment of approximately 30 active neutral accounts per day on October 11. Left-leaning, right-leaning, and balanced accounts began appearing consistently in the dataset around October 7, with each group reaching a stable count of about 30 active accounts per day shortly thereafter. Each neutral account receives approximately 500 tweets per session, while each left-leaning, right-leaning, and balanced account receives around 700 tweets per session.
[…]
Number of bots
[…]
Number of tweets
Technical Considerations for Neutral Accounts.
Managing neutral accounts presented several challenges during data collection. For accounts that followed no users, $\mathbb{X}$ disabled the timeline after 7 days, requiring us to create additional neutral bots to maintain at least 30 active accounts daily. However, during the election period, the platform temporarily modified this restriction, disabling the timeline immediately after account creation for accounts that followed no users. Consequently, data collection was limited to approximately 10 older neutral accounts after November 5.
Appendix B Top Recommended Users in Left-Leaning, Right-Leaning, and Balanced Accounts
Figure 6, figure 7 and figure 8 display the top 20 recommended users in left-leaning, right-leaning, and balanced accounts, ranked by their average weighted occurrence per 1,000 tweets. Each box in the boxplot represents the distribution of exposure across all accounts in each group, with red indicating right-leaning users and blue indicating left-leaning users. Political leanings of users are inferred based on publicly available data, which may be subject to inaccuracies or changes over time. Notably, balanced accounts receive a roughly even mix of left- and right-leaning recommendations, whereas left- and right-leaning accounts predominantly receive recommendations from ideologically aligned users.
@@ Appendix B Top Recommended Users in Left-Leaning, Right-Leaning, and Balanced Accounts
Appendix C Profile Information of the Top Recommended Users
@@ 11 AMAZlNGNATURE Nature is Amazing Animals Nature Discovery 4496570
21 ProjectLincoln The Lincoln Project “You cannot escape the responsibility of tomorrow by evading it today.” – Abraham Lincoln — Home of #TheBreakdown and LP Podcast 2994780
@@ 28 DC_Draino DC_Draino Rogan O’Handley 1855070
32 PopBase Pop Base Pop Base is your best source for all pop culture related entertainment, news, award show coverage, chart updates, statistics and more. — email@popbase.tv 1683990
33 CollinRugg Collin Rugg Co-Owner of Trending Politics — Investor — American 1561596
@@
40 Riley_Gaines_ Riley Gaines Host of Gaines for Girls podcast — Author of Swimming Against the current — TPUSA contributor — Director of the Riley Gaines Center 1283591
@@ 49 Dexerto Dexerto The leading source for influencer, streamer, gaming, and viral content 980351
56 Angry_Staffer Angry Staffer Not a WH Staffer — Politics, NatSec, and Snark - Your Mileage May Vary — Subscribe to my Patreon newsletter for free: https://t.co/Kj4zTIcPyk — 609103
@@
60 acnewsitics Alex Cole Software Engineer & Pilot — Progressive Follow @newsitics & https://t.co/Retehye9rD 286475
@@ 61 Logically_JC John Collins Dad Husband Low-Key Nerd EdD / JD 225758
66 canlesofficial Canles Engineered for walking — Comfy & versatile footwear for life’s adventures — Breathable, lightweight designs 7407